OpenStudy (anonymous):

When a certain number divides 67,82,97, they all leave a remainder of 7. Find the largest possible value of the number.

OpenStudy (asnaseer):

lets call the unknown number x. then you can write this as:\[67=ax+7\]\[82=bx+7\]\[97=cx+7\]giving us:\[ax=60, bx=75, cx=90\]where a, b and c are some other unknown integers. so now you just need to find the GCd (Greatest Common Divisor) of 60, 75 and 90

OpenStudy (anonymous):

so confusing?

OpenStudy (asnaseer):

which part do you find difficult to understand?

OpenStudy (anonymous):

OOO nvm

OpenStudy (anonymous):

i get it

OpenStudy (asnaseer):

ok - good

OpenStudy (asnaseer):

do you know how to find GCD of the 3 numbers?

OpenStudy (anonymous):

No

OpenStudy (asnaseer):

write them out as a product of their primes factors - that should help

OpenStudy (anonymous):

15?

OpenStudy (asnaseer):

\[60=2^2*3*5\]\[75=3*5^2\]\[90=2*3^2*5\]

OpenStudy (asnaseer):

then try and spot what is COMMON between all of them

OpenStudy (asnaseer):

e.g. 2 is not comon as it is not a factor of 75

OpenStudy (asnaseer):

3 IS common - can you spot another common factor?

OpenStudy (anonymous):

5?

OpenStudy (asnaseer):

good - so 3 and 5 are the only common factors. so the answer is x=3*5=15

OpenStudy (anonymous):

thank you

OpenStudy (asnaseer):

yw